OpenStudy (anonymous):

Can someone explain: Simplify 2 times the square root of 108 over the square root of 180y???

OpenStudy (anonymous):

Looks like this? \[\frac{2\sqrt{108}}{\sqrt{180y}}\]

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Okay. Basically, the problem is asking you to 1) remove all the square roots from the denominator and 2) make sure there are no perfect squares in the radical.

OpenStudy (anonymous):

The first thing you're going to do is multiply the top and bottom by 1, but in sneaky way. Remember that anything divided by itself is 1, so \[1=\frac{\sqrt{180y}}{\sqrt{180y}}\] Correct?

OpenStudy (anonymous):

Correct! :)